# Indoor Location Fingerprinting Privacy: A Comprehensive Survey

### v1

@@ Abstract.
The pervasive integration of Indoor Positioning Systems (IPS) arises from the limitations of Global Navigation Satellite Systems (GNSS) in indoor environments, leading to the widespread adoption of Location-Based Services (LBS). Specifically, indoor location fingerprinting employs diverse signal fingerprints from user devices, enabling precise location identification by Location Service Providers (LSP). Despite its broad applications across various domains, indoor location fingerprinting introduces a notable privacy risk, as both LSP and potential adversaries inherently have access to this sensitive information, compromising users’ privacy. Consequently, concerns regarding privacy vulnerabilities in this context necessitate a focused exploration of privacy-preserving mechanisms. In response to these concerns, this survey presents a comprehensive review of Privacy-Preserving Mechanisms in Indoor Location Fingerprinting (ILFPPM) based on cryptographic, anonymization, differential privacy (DP), and federated learning (FL) techniques. We also propose a distinctive and novel grouping of privacy vulnerabilities, adversary and attack models, and available evaluation metrics specific to indoor location fingerprinting systems. Given the identified limitations and research gaps in this survey, we highlight numerous prospective opportunities for future investigation, aiming to motivate researchers interested in advancing this field. This survey serves as a valuable reference for researchers and provides a clear overview for those beyond this specific research domain.
Indoor Localization, Privacy-Preserving, Location Fingerprinting
ccs: Security and privacy Privacy protections
1. Introduction
@@
Localization is the process of determining the spatial coordinates or position of a tracked item with respect to several reference points within a predefined area (Geng et al., 2021). There are growing appeals for providing Location-based services (LBS) to people around the world. The appearance of Global Navigation Satellite Systems (GNSS) was a turning point in localization, but lately, the traditional scope of positioning systems evolved with the advent of mobile computing, including the Internet of Things (IoT) and wearable devices, moving from military tracking and civilian navigation to location information. For both personal and commercial purposes, location data is now a crucial component in bridging the gap between the real and digital worlds. Nowadays, due to the widespread use of mobile devices, users are constantly generating various location data during their daily activities. Accordingly, there is a growing interest in LBS, which offer services depending on the user’s location. Such services assist people in finding the appropriate route to different destinations, selecting various points of interest (POI) based on their preferences, and even remembering to take precautions for their healthcare purposes (Pascacio et al., 2021).
@@
All these LBSs require reliable and real-time location information, which is often accomplished using GNSS technologies. However, due to the severe attenuation of satellite signals and the low visibility of satellites in indoor environments, they mostly fail for indoor localization purposes (Asaad and Maghdid, 2022). The main indoor localization technologies are alternative signals such as Wi-Fi, Bluetooth, FM, AM, GSM or LTE (Moghtadaiee and Dempster, 2014; Rahman et al., 2017). During localization, they all should face indoor area challenges such as non-line-of-sight (NLOS) problems, multipath propagation, signal blockage, intentional and unintentional interferences, etc.
@@
Indoor positioning has gotten so much attention in the last decade as individuals spend 80% of their time indoors (Kepeis et al., 2001), which shows the significant role of indoor localization. Unlike in outdoor positioning systems, the user has no direct access to her location, and it is calculated by the Location Service Provider (LSP), the server in the Indoor Positioning System (IPS), based on the signal information received from the user, as shown in Fig. 1. Moreover, more accuracy is required for indoor localization than it is for outdoor localization which only needs meter-level accuracy. The indoor environments are diverse and include residences, workplaces, shopping malls, warehouses, hospitals, and retail centers. Additionally, there are many various applications for end users with different requirements for accuracy and coverage. For instance, while law-enforcement applications with urban/rural coverage have an accuracy of a few meters, ambient assisted-living applications require room-level coverage with an accuracy below $1m$ (Mautz, 2012). Therefore, no IPS has evolved as an all-encompassing solution due to the unique characteristics of indoor settings and the diversity of applications. Not only the coverage and accuracy are important to consider when selecting the base positioning technology, but also the deployment and maintenance costs are crucial. The diverse solutions are a clear indicator that there is no single alternative for GNSS in indoor areas, and different positioning technologies co-exist.
Figure 1. Outdoor vs. indoor positioning systems.
However, with all the benefits provided by IPS, users are deeply concerned about their privacy when their personal location data is accessed; hence, proper privacy preservation is of utmost importance for indoor LBS applications, motivated many researchers to work on how to protect the privacy of user locations in indoor environments. The development of localization methodologies has also made it possible for indoor LBS providers to quickly gather a significant amount of location data from several users, which is necessary to better understand user behavior and, as a result, improve the quality of the services, but it severely violates the privacy of users at the same time.
In response to privacy concerns, many countries have enacted privacy laws and regulations, such as the General Data Protection Regulation (GDPR) [^1] https://gdpr-info.eu/ in the European Union, and the California Consumer Privacy Act (CCPA) [^2] https://oag.ca.gov/privacy/ccpa in the United States, to provide individuals with greater control over their personal data, including location information. Specifically for location privacy, an architecture of an indoor localization model ensuring GDPR compliance through the integration of a specialized GDPR-based access control system is introduced in (Barsocchi et al., 2020).
@@
Therefore, the main obstacle to the widespread adoption of IPS in the world is privacy. When a user is linked to the network, IPS always ”knows” where they are. Without the user’s prior consent, tracking her location is prohibited as users are generally reluctant to share their whereabouts. Location data reveal private and sensitive information on each user, including their health conditions, their interests, their views and behaviors, their usage amount of electricity, their workplaces, and their home and job locations. The leakage of private assembly line information, sensitive process information, and other relevant information are additional security risks posed by IPS in commercial or high-profile buildings. Moreover, the radio map of the buildings can be used to access the structural configuration, which might raise terrorist attacks and other negative actions. Therefore, practical answers are required to these privacy problems (Singh et al., 2021). Authorities will establish stringent but pertinent norms and regulations for industries to use users’ location data as nontechnical solutions. Without the user’s explicit consent, data should not be shared with third-party organizations or people. Therefore, the lack of comprehensive efforts on this topic motivates us to examine indoor location privacy from various points of view thoroughly.
@@
Several papers are surveying IPS in terms of accuracy, localization techniques, and technologies (Singh et al., 2021; Nessa et al., 2020; Zafari et al., 2019; Ngamakeur et al., 2020). Considering the security and privacy of IPS, on the other hand, there are two limited surveys (Holcer et al., 2020; YILDIZ and PEKEY, 2019) and only one comprehensive survey (Sartayeva and Chan, 2023).
The first review paper in 2020, (Holcer et al., 2020), conducts a limited analysis of privacy in IPS using the Preferred Reporting Items for Systematic Review and Meta-Analysis (PRISMA) guidelines (Liberati et al., 2009). The paper categorizes the existing research papers (up to 2020) into three groups: privacy on the device, during transmission, and at the server, and explains them based on this categorization. However, as this research is from 2020, the research studies after that time are not included. In 2022, the only comprehensive survey (Sartayeva and Chan, 2023) categorizes the indoor positioning methods into two items: (1) non-collaborative methods including proximity-based methods, geometry methods, fingerprinting, and others, and (2) collaborative methods including mobile proximity-based methods, mobile geometric methods, and others. For the sake of privacy and security, the survey applies a confidentiality, integrity, and authenticity trilogy to each of the collaborative and non-collaborative methods. The main focus of this paper, however, is on the security issues of IPS. Regarding various privacy attacks and privacy-preserving techniques when location fingerprinting technique is used, it only mentions limited number of publications. The last existing survey in 2023, (YILDIZ and PEKEY, 2019), first investigates the existing reviews on IPS and the technologies and techniques applied to them, followed by their advantages and disadvantages. It explains some of the major challenges for IPS. In terms of privacy in indoor location fingerprinting, it surveys a restricted number of studies, 15 papers between 2018 and 2022, without employing categorization, most of which are based on secure two-party computation and homomorphic encryption. Furthermore, it neither takes into account more recent papers introducing new methodologies nor discusses privacy attacks.
@@
This paper focuses on privacy-preserving mechanisms in the context of indoor location fingerprinting, the most widely used localization technique in IPS. Throughout the rest of the paper, when referring to IPS, we specifically mean IPS utilizing location fingerprinting as its localization technique. The contributions of this survey paper are outlined as follows:
@@
  • To the best of our knowledge, this is the first deep dive survey considering indoor location fingerprinting privacy-preserving mechanisms (ILFPPM) from various novel aspects: indoor location privacy definitions, applications, entities vulnerabilities, possible attacks, privacy protection techniques, challenges, privacy metrics, and vision of future research directions.
@@
  • In addition to discussing IPS fundamentals, our exploration includes an thorough analysis of the various applications of IPS, investigating their practical implementations and addressing the privacy concerns associated with the collection, storage, and utilization of location data in indoor spaces.
@@
  • We comprehensively identify privacy vulnerabilities and all sources of privacy leakages in indoor location fingerprinting systems, examining them from diverse perspectives encompassing system entities, indoor location data structure, and information inference. Additionally, we introduce a novel categorization of adversary models and attack models specific to these systems, marking the first systematic classification in this domain
@@
  • To facilitate empirical investigations into ILFPPM, we provide list of datasets from prior research. We also categorize, for the first time, all metrics employed in existing studies. This endeavor is intended to establish a robust benchmark within the research community, fostering future empirical analyses and enhancing technical insights into this domain.
@@
  • Through the literature analysis, we have outlined the unresolved challenges and suggested various potential paths for future research in location privacy attacks and ILFPPM. Our aim is to motivate readers with an interest in this field to further explore these issues.
@@
As outlined in Fig. 2, the rest of the paper is organized as follows. Section 2 presents IPS fundamentals including methodology, techniques, and technologies. Section 3 explains applications of IPS with their privacy concerns. Section 4 discusses the sources of privacy leakages in indoor location fingerprinting, followed by the adversary and attack models in Section 5. The existing ILFPPM are fully reported in Section 6. All the datasets and metrics utilized in these studies also presented in this Section. Section 7 discusses the future research directions, and Section 8 concludes the paper.
Figure 2. Taxonomy of this paper.
2. Indoor Positioning System Fundamentals
@@
To understand privacy protection methods, it is important to first understand the fundamentals of IPS. IPS is a GPS-free system that continuously predicts the positions of people or objects in indoor environments by first applying a distance estimation algorithm, followed by the localization algorithm (Liu et al., 2007). Using a suitable ranging technique, an IPS calculates the distance between the target and anchor nodes whose coordinates are known a priori during the distance measurement phase. Using these distance observations, the IPS then uses several localization/positioning techniques to approximate the target’s location.
2.1. Ranging Techniques
@@
Here we discuss various ranging techniques that have been employed for indoor localization so far.
2.1.1. Received Signal Strength (RSS)
@@
In general, Received Signal Strength (RSS) represents the intensity of a signal measured at the receiver, typically at a certain distance from a signal source. It serves as one of the most straightforward signal metrics to measure, providing valuable insights into the strength of wireless signals. However, particularly in indoor environments, RSS measurements can be susceptible to various factors that introduce inaccuracies. These factors include fading, shadowing, refraction, scattering, and reflections within complex indoor layouts.
Due to the sensitivity of RSS measurements to these environmental influences, they often yield erroneous distance estimates in IPS. To mitigate the fluctuations and inaccuracies associated with RSS, various techniques and machine learning (ML) approaches have been adopted (Ben Kilani et al., 2014). These methods aim to enhance the reliability and precision of RSS-based positioning. Some of the strategies employed include the implementation of sophisticated filters and the utilization of ML algorithms. These approaches are designed to filter out noise, account for signal variations, and extract meaningful information from RSS data, ultimately improving the performance and robustness of IPS.
@@
When measuring the RSS fingerprinting values is not feasible in a specific environment, or a very emergency training procedure is required, one way is to generate fingerprinting RSS values instead. For generating RSS values in an indoor area, the long-distance path-loss model has been frequently utilized, as it demonstrates how RSS measurements vary with distance from an AP (Chen and Terzis, 2011). The RSS values sensed from the $j$’th AP, ${S}_{u,j}$(dBm) at a physical distance $d_{j}$ in meter, are calculated as follows:
(1) ${RSS}_{u,j}={s_{d_{0}}}_{j}-10\alpha_{j}\log\left(\frac{d_{j}}{d_{0}}\right)+X_{\sigma_{j}},$
@@
where ${s_{d_{0}}}_{j}$dBm is the power received at $d_{0}$m (usually $1$m) from $j$th AP, and $\alpha_{j}$ is the path-loss exponent associated with that AP. Due to the shadowing in indoor environments, $X_{\sigma_{j}}\sim\mathcal{N}(0,\sigma_{j})$ models path-loss variation at a single point and is assumed to be a zero mean Gaussian random variable with a standard deviation given by $\sigma_{j}$dB.
2.1.2. Signal Propagation Time

### v2

# A Survey and Future Outlook on Indoor Location Fingerprinting Privacy Preservation

@@ Abstract
The pervasive integration of Indoor Positioning Systems (IPS) arises from the limitations of Global Navigation Satellite Systems (GNSS) in indoor environments, leading to the widespread adoption of Location-Based Services (LBS) in places such as shopping malls, airports, hospitals, museums, corporate campuses, and smart buildings. Specifically, indoor location fingerprinting (ILF) systems employ diverse signal fingerprints from user devices, enabling precise location identification by Location Service Providers (LSP). Despite its broad applications across various domains, ILF introduces a notable privacy risk, as both LSP and potential adversaries inherently have access to this sensitive information, compromising users’ privacy. Consequently, concerns regarding privacy vulnerabilities in this context necessitate a focused exploration of privacy-preserving mechanisms. In response to these concerns, this survey presents a comprehensive review of Indoor Location Fingerprinting Privacy-Preserving Mechanisms (ILFPPM) based on cryptographic, anonymization, differential privacy (DP), and federated learning (FL) techniques. We also propose a distinctive and novel grouping of privacy vulnerabilities, adversary models, privacy attacks, and evaluation metrics specific to ILF systems. Given the identified limitations and research gaps in this survey, we highlight numerous prospective opportunities for future investigation, aiming to motivate researchers interested in advancing ILF systems. This survey constitutes a valuable reference for researchers and provides a clear overview for those beyond this specific research domain. To further help the researchers, we have created an online resource repository, which can be found at https://github.com/amir-ftlz/ilfppm.
keywords:
Indoor localization , Location privacy , Privacy-preserving , Location fingerprinting
1 Introduction
@@
Localization is the process of determining the spatial coordinates or position of a tracked item with respect to several reference points within a predefined area [1]. There is an increasing demand for offering Location-Based Services (LBS) to individuals worldwide. The appearance of Global Navigation Satellite Systems (GNSS) was a turning point in localization, but lately, the traditional scope of positioning systems evolved with the advent of mobile computing, including the Internet of Things (IoT) and wearable devices, moving from military tracking and civilian navigation to location information. Location data has become essential in connecting the real and digital worlds for both personal and commercial uses. Nowadays, due to the widespread use of mobile devices, users are constantly generating various location data during their daily activities. Accordingly, there is a growing interest in LBS, which offers services depending on the user’s location. Such services assist people in finding the appropriate route to different destinations, selecting various points of interest (POI) based on their preferences, and even remembering to take precautions for their healthcare purposes [2].
@@
All these LBSs require reliable and real-time location information, which is often accomplished using GNSS technologies. However, due to the severe attenuation of satellite signals and the low visibility of satellites in indoor environments, they mostly fail for indoor localization purposes [3]. The main indoor localization technologies are alternative signals such as Wi-Fi, Bluetooth, FM, AM, GSM, or LTE [4, 5]. During localization, they should face indoor area challenges such as non-line-of-sight (NLOS) problems, multipath propagation, signal blockage, intentional and unintentional interferences, etc.
@@
Indoor positioning has gotten so much attention in the last decade as individuals spend 80% of their time indoors [6], which shows the significant role of indoor localization. Unlike in outdoor positioning systems, the user has no direct access to her location, and it is calculated by the Location Service Provider (LSP), the server in the Indoor Positioning System (IPS), based on the signal information received from the user, as shown in Fig. 1. Additionally, indoor localization demands higher precision compared to outdoor localization, which typically only requires meter-level accuracy. The indoor environments are diverse, including residences, workplaces, shopping malls, warehouses, hospitals, and retail centers. Additionally, there are various applications for end users with different requirements for accuracy and coverage. For instance, while law-enforcement applications with urban/rural coverage have an accuracy of a few meters, ambient assisted-living applications require room-level coverage with an accuracy below $1m$ [7]. Therefore, no IPS has evolved as an all-encompassing solution due to the unique characteristics of indoor settings and the diversity of applications. The coverage and accuracy are important to consider when selecting the base positioning technology, and the deployment and maintenance costs are crucial. The diverse solutions are a clear indicator that there is no single alternative for GNSS in indoor areas and different positioning technologies co-exist.
Figure 1: Outdoor vs. indoor positioning systems (no privacy is considered).
@@
Therefore, one of the main obstacles to the widespread adoption of IPS in the world is privacy. When a user is linked to the network, IPS always ”knows” where they are. Without the user’s prior consent, tracking her location is prohibited as users are generally reluctant to share their whereabouts. Location data reveal private and sensitive information on each user, including their health conditions, their interests, their views and behaviors, their usage amount of electricity, their workplaces, and their home and job locations. The leakage of private assembly line information, sensitive process information, and other relevant information are additional security risks posed by IPS in commercial or high-profile buildings. Moreover, the radio map of the buildings can be used to access the structural configuration, which might raise terrorist attacks and other negative actions. Therefore, practical answers are required to these privacy problems [8]. Authorities will establish stringent but pertinent norms and regulations for industries to use users’ location data as nontechnical solutions. Without the user’s explicit consent, data should not be shared with third-party organizations or people. Therefore, the lack of comprehensive efforts on this topic motivates us to examine indoor location privacy from various points of view thoroughly.
1.1 Existing Surveys and Our Contributions
@@
Several papers are surveying IPS in terms of accuracy, localization techniques, and technologies [8, 9, 10, 11]. Considering the security and privacy of IPS, on the other hand, there are two limited surveys [12, 13] and only one comprehensive survey [14]. The comparison of these three papers with our paper is reported in Table 1, using labels such as Y (Yes), N (No), and L (Limited) to indicate the extent to which each survey covers specific subject.
The first review paper in 2020, [12], conducts a limited analysis of privacy in IPS using the Preferred Reporting Items for Systematic Review and Meta-Analysis (PRISMA) guidelines [15]. The paper categorizes existing research papers (up to 2020) into three main groups: privacy on the device, during transmission, and at the server. The authors identify several privacy challenges specific to each category, such as the risk of unauthorized data collection on user devices, the interception of sensitive data during wireless transmission, and the vulnerabilities associated with centralized storage solutions, which could become targets for data breaches or misuse. However, as this research was conducted in 2020, it does not account for advancements in privacy-preserving techniques such as federated learning (FL), differential privacy (DP), or new cryptographic protocols that have since emerged.
The other survey in 2023 [13] starts by examining existing reviews on IPS and analyzes the technologies and techniques applied to them while discussing their advantages and disadvantages. This paper focuses on recent developments and surveys of a restricted number of studies— 15 papers published between 2018 and 2022— mostly related to privacy-preserving techniques in indoor location fingerprinting (ILF), most of which utilize secure two-party computation and homomorphic encryption. Although these cryptographic techniques are discussed in depth, the survey lacks a broader categorization of privacy threats and attacks, focusing primarily on encryption-based solutions without much emphasis on alternative methods such as anonymization, DP, or FL. Moreover, the paper does not take into account more recent research papers introducing novel privacy-preserving methods. Additionally, it does not discuss the potential vulnerabilities of these cryptographic techniques, leaving out a discussion on their performance trade-offs, scalability, or potential exposure to specific privacy attacks. Despite covering some technical challenges for IPS, this survey misses the opportunity to address privacy in a broader, evolving landscape of IPS technologies.
Also in 2023, a comprehensive survey [14] provides a broad overview of security and privacy issues in indoor positioning, encompassing a range of methods such as proximity-based, geometric, and collaborative techniques. It categorizes indoor positioning methods into two major categories: (1) non-collaborative methods, which include proximity-based methods, geometry methods, location fingerprinting, and others, and (2) collaborative methods, which include mobile proximity-based methods, mobile geometric methods, and others. For the sake of privacy and security, the survey applies the confidentiality, integrity, and authenticity (CIA) trilogy to both collaborative and non-collaborative methods. While their work presents a valuable analysis of security threats (e.g., jamming, replay attacks, data tampering), our paper takes a specialized approach by focusing exclusively on ILF privacy, an area of growing importance as ILF becomes more prevalent in real-world applications. One of the main differences in our work is the identification and detailed analysis of privacy-vulnerable entities specific to ILF systems. Unlike the broader scope in [14], which touches lightly on privacy concerns, our survey breaks down the unique risks each of ILF systems’ entities faces, offering a more granular examination of how privacy can be compromised in fingerprinting-based systems. Additionally, our paper offers a thorough discussion on location privacy attacks specifically targeting ILF through signal fingerprints. These attacks, which exploit the inherent characteristics of ILF data, pose distinct threats that are not fully explored in these three existing surveys. In terms of solutions, while authors in [14] focus on general security and privacy mechanisms, our paper thoroughly introduces advanced privacy-preserving mechanisms specifically for ILF. These include various cryptographic approaches, anonymization, DP, and FL, each critically evaluated for its effectiveness in ILF contexts. We also address the practical challenge of balancing accuracy and privacy in these systems— an essential consideration for real-world deployments, where maintaining system utility without sacrificing privacy is crucial.
Table 1: Comparison of existing surveys on indoor localization privacy (Y: Yes, N: No, L: Limited)
Subject [12] [13] [14] Our survey
ILF Focus N Y N Y
Applications of IPS N Y Y Y
ILF Entities’ Privacy Vulnerabilities N N N Y
Adversary Models N N Y Y
Categorization of Privacy Attacks N N N Y
Privacy Metrics N N N Y
Datasets N N N Y
Cryptographic Techniques for ILF L Y L Y
Anonymization Techniques for ILF N N L Y
Differential Privacy (DP) for ILF N L L Y
Federated Learning (FL) for ILF N N N Y
Comparative Analysis of Privacy Techniques N N N Y
Case Studies on Privacy Violations in ILF N N N Y
Challenges and Future Directions L N Y Y
@@
Our paper, as mentioned, focuses on privacy-preserving mechanisms in the context of ILF, the most widely used localization technique in IPS. Incorporating the latest research and newly developed methods for addressing privacy concerns, this survey provides a comprehensive review of existing privacy-preserving solutions for ILF systems, including an in-depth analysis of privacy involving the entities in the IPS, the metrics and datasets commonly used, and the various attacks these systems face. Note that throughout the rest of the paper, when referring to IPS, we specifically mean IPS utilizing location fingerprinting as its localization technique. The main contributions of this survey paper are outlined as follows:
@@
  • To the best of our knowledge, this is the first deep dive survey considering indoor location fingerprinting privacy-preserving mechanisms (ILFPPM) from various novel aspects: indoor location privacy definitions, applications, entities vulnerabilities, possible attacks, privacy metrics, privacy protection techniques, challenges, and vision of future research directions.
@@
  • In addition to discussing IPS fundamentals, our exploration includes a thorough analysis of the various applications of IPS, investigating their practical implementations, and addressing the privacy concerns associated with the collection, storage, and utilization of location data in indoor spaces.
@@
  • We comprehensively identify privacy vulnerabilities and all sources of privacy leakages in ILF systems for the first time. We also examine them from diverse perspectives encompassing system entities, indoor location data structure, and information inference. Additionally, we introduce a novel categorization of adversary models and privacy attack models specific to these systems, marking the first systematic classification regarding privacy in ILF systems.
@@
  • To facilitate empirical investigations into ILFPPM, we provide a list of datasets from prior research. We also categorize all metrics (regarding localization, quality of service (QoS), and privacy) employed in existing studies for the first time, which is intended to establish a robust benchmark within the research community, fostering future empirical analyses and enhancing technical insights into ILF systems.
@@
  • Through the literature analysis, we have outlined the unresolved challenges and suggested various potential paths for future research in ILF privacy attacks and ILFPPM.
@@
As outlined in Fig. 2, the rest of the paper is organized as follows. Section 2 presents IPS fundamentals including methodology, techniques, and technologies. Section 3 discusses the sources of privacy leakages in ILF, followed by the adversary and attack models in Section 4. The existing ILFPPM are fully reported in Section 5. All the datasets and metrics utilized in these studies are also presented in this Section. Section 6 explains applications of IPS with their privacy concerns. Section 7 discusses the future research directions, and Section 8 concludes the paper. The abbreviations used in this paper are also listed in Table 2.
Figure 2: Taxonomy of the paper.
2 Indoor Positioning System Fundamentals
@@
Understanding the fundamentals of IPS is essential before delving into privacy protection methods. IPS is a GPS-free system that continuously predicts the positions of people or objects in indoor environments by initially applying a distance estimation algorithm, followed by the localization algorithm [16]. Using a suitable ranging technique, an IPS calculates the distance between the target and anchor nodes whose coordinates are known a priori during the distance measurement phase. Using these distance observations, the IPS then uses several localization/positioning techniques to approximate the target’s location.
Table 2: Summary of important abbreviations.
Notation Description Notation Description
IPS Indoor Positioning System ILF Indoor Location Fingerprinting
LPPM Location Privacy Preserving Mechanism ILFPPM Indoor Location Fingerprinting Privacy Preserving Mechanism
LBS Location Based Service LSP Location Service Provider
RP Reference Point AP Access Point
RSS Received Signal Strength CSI Channel State Information
AoA Angle of Arrival PoA Phase of Arrival
ToA Time of Arrival TDoA Time Difference of Arrival
PoI Point of Interest IoT Internet of Things
ML Machine Learning NLOS Non-Line Of Sight
DP Differential Privacy CDP Centralized Differential Privacy
LDP Local Differential Privacy FL Federated Learning
MPC Multi-Party Computation 2PC Two-Party Computation
2.1 Ranging Techniques
@@
In this section, we explore the various ranging techniques that have been used for indoor localization up to this point.
2.1.1 Received Signal Strength (RSS)
@@
In general, Received Signal Strength (RSS) represents the intensity of a signal measured at the receiver, typically at a certain distance from a signal source. It is one of the most straightforward signal metrics to measure, providing valuable insights into the strength of wireless signals. However, particularly in indoor environments, RSS measurements can be susceptible to various factors that introduce inaccuracies. These factors include fading, shadowing, refraction, scattering, and reflections within complex indoor layouts.
@@
The long-distance path-loss model has been frequently utilized for generating RSS values in an indoor area, as it demonstrates how RSS measurements vary with distance from an AP [17]. The RSS values sensed from the $j$’th AP, ${S}_{u,j}$(dBm) at a physical distance $d_{j}$ in meter, are calculated as follows:
${RSS}_{u,j}={s_{d_{0}}}_{j}-10\alpha_{j}\log\left(\frac{d_{j}}{d_{0}}\right)+X_{\sigma_{j}},$ (1)
@@
where ${s_{d_{0}}}_{j}$dBm is the power received at $d_{0}$m (usually 11m) from $j$th AP, and $\alpha_{j}$ is the path-loss exponent associated with that AP. Due to the shadowing in indoor environments, $X_{\sigma_{j}}\sim\mathcal{N}(0,\sigma_{j})$ models path-loss variation at a single point and is assumed to be a zero mean Gaussian random variable with a standard deviation given by $\sigma_{j}$dB.
Due to the sensitivity of RSS measurements to these environmental influences, they often yield erroneous distance estimates in IPS. To mitigate the fluctuations and inaccuracies associated with RSS, various techniques and machine learning (ML) approaches have been adopted [18]. These methods aim to enhance the reliability and precision of RSS-based positioning. Some of the strategies employed include the implementation of sophisticated filters and the utilization of ML algorithms. These approaches are designed to filter out noise, account for signal variations, and extract meaningful information from RSS data, ultimately improving the performance and robustness of IPS.
2.1.2 Signal Propagation Time